OpenStudy (anonymous):

If 3a=2b and 3b=5c, what is the ratio of a to c?

OpenStudy (anonymous):

First you should try to set each equation equal to the same thing. Since you are trying to find the ratio of a to c, you should set both equal to b.

OpenStudy (anonymous):

3a=2b b=(3a)/2 3b=5c b=(5c)/3

OpenStudy (anonymous):

k

OpenStudy (anonymous):

since b=(3a)/2 and b=(5c)/3, (5c)/3=(3a)/2. Now you can simplify and get the ratio.

OpenStudy (anonymous):

so

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

5/3=1.67

OpenStudy (anonymous):

3/2=1.5

OpenStudy (anonymous):

a=1.5

OpenStudy (anonymous):

c=1.67

OpenStudy (anonymous):

1.67/1.5=10/9

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

that's the answer in the book...

OpenStudy (anonymous):

then yeah, I guess that's it haha. I thought you were going for an answer more like 9a=10c

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ty